# Gated KalmaNet: A Fading Memory Layer Through Test-Time Ridge Regression

### v1

@@ Abstract
As efficient alternatives to softmax Attention, linear state-space models (SSMs) achieve constant memory and linear compute, but maintain only a lossy, fading summary of the past, often leading to inferior performance in recall oriented tasks. We propose Gated KalmaNet (GKA), a layer that reduces this gap by accounting for the full past when predicting the next token, while maintaining SSM-style efficiency. GKA achieves this by solving an online ridge regression problem at test time, with constant memory and linear compute cost in the sequence length. Drawing inspiration from the Kalman Filter, we iteratively solve the online ridge regression problem. However, a critical insight is that standard Kalman filter equations are numerically unstable in low-precision environments (like bfloat16) and difficult to parallelize in modern hardware. We address both challenges through two key innovations: (1) an adaptive regularization strategy with input-dependent gating that controls the condition number of the ridge regression, ensuring numerical stability while balancing memory retention. And (2) the use of Chebyshev Iteration instead of other conventional iterative solvers, which we demonstrate to be more stable in low-precision settings. To further improve scalability, we develop a hardware-aware chunk-wise implementation of Chebyshev Iteration along with custom kernels for backpropagating through our adaptive regularization and gating mechanisms. Empirically, GKA shows strong language understanding capabilites on short-context tasks outperforming existing SSM layers (like Mamba2, GLA and Gated DeltaNet). On long-context, GKA excels at real-world RAG and LongQA tasks up to 128k tokens, achieving more than 1010% relative improvement over other fading memory baselines.
@@ 2 Prior Work and Preliminaries
Similarly to that of Eq.˜Attn the case with low-rank-plus-identity matrices can often be justified from an optimization perspective. For example, Gated DeltaNet [schlag2021linear, Yang-NeurIPS2024] updates the state via (II is the $D\times D$ identity matrix)
@@
which can be viewed as applying one gradient descent step with stepsize $\beta_{t}$ and initialization $\gamma_{t}S_{t-1}$ to the objective
@@ 3.1 Motivation from Kalman Filter
  • In Eq.˜1, the weights of the same residual vary over time as the queries differ at each time, while in Eq.˜3 $i$-th weight $\eta_{i}$ is constant once observed at time ii. The former results in quadratically many weights—thus a quadratic time complexity—and the latter linearly many.
@@ 4.1 CH with Adaptive Regularization & Weighting
With the weighted covariances $U_{t}:=\sum_{i=1}^{t}\eta_{i}v_{i}k_{i}^{\top}$ and $H_{t}:=\sum_{i=1}^{t}\eta_{i}k_{i}k_{i}^{\top}$, we note that $y_{t}$ can be computed via first solving $(H_{t}+\lambda I)x=q_{t}$ for $x$ and then left-multiplying $U_{t}$. An exact solver (e.g., torch.linalg.solve) can do so with high accuracy, by parallelizing over the batch dimension. However, it is inefficient here for two reasons. First, it takes $O(D^{3})$ time for every $t$. Second, it requires explicitly forming and materializing all $H_{t}$’s, which would entail a large I/O cost. In light of this, we resort to first-order iterative methods that admit chunk-wise implementation without materializing all $H_{t}$’s, enabling parallelism over chunks and batches. Furthermore, they often take $O(D^{2})$ time complexity per iteration and can converge quickly in a few iterations. The iterative method we choose is the Chebyshev Iteration (CH); we proceed to describe its basic idea, with a justification of using CH deferred to Section˜4.2.3.
@@ 4.2.2 Backward Pass
Chunk-wise Recurrence. As indicated, a chunk-wise implementation for computing $A_{i}k_{i}$ is known. On the other hand, computing $B_{i}k_{i}$ is more challenging than $A_{i}k_{i}$, as the additive term $w_{i}H_{i}$ in the backward recursion Eq.˜9 is not necessarily rank-11 ; rather, $H_{i}$ itself is defined via the forward recursion in Eq.˜CH. Our contribution here is a derivation for computing $B_{i}k_{i}$ efficiently in a chunk-wise manner.
@@ 4.2.3 Comparison to Other Iterative Solvers
In Fig.˜1b, CG (impl) as a standalone layer has its gradient close to that of the exact solver up to a $10^{-3}$ relative difference. In Fig.˜1c, this difference is amplified to almost 11 in a 5-layer LLAMA where Eq.˜Attn is replaced with Eq.˜3. This indicates CG (impl) completely deviates from the reference gradient (exact), defeating its purpose of training the network from the regression feedback. In contrast, the gradients of CH (impl) and CH (auto) are eventually close to that of the exact solver either as a single layer (Fig.˜1c) or within multiple layers (Fig.˜1d), up to a $10^{-6}$ difference. Moreover, the curves for CH (impl) and CH (auto) nearly overlap, suggesting that their gradients may be close. The following lemma confirms and formalizes this intuition (see Appendix˜C for a proof), thereby justifying our choice of CH over the alternatives:
@@ 5.1 GKA on Synthetic Associative Recall Tasks
We compare GKA with Attention and other linear SSM baselines on this task. For each memory layer type, we train 2-layer models on MQAR training data and evaluate on a held-out test set. We repeat this experiment for 44 different learning rates spanning from $10^{-4}$ to $10^{-2}$. As shown in Fig.˜3a, GKA improves upon every other linear SSM baseline at all sequence lengths and model dimensions considered. Note, the complexity of the task increases with increasing sequence length and number of key-value pairs, while larger model dimensions improve memorization capacity through increased state size. The success of our layer can be attributed to our modeling choice: unlike other fading memory designs (like GDN or Mamba2), we construct states based on the optimal MAP estimate conditioned on the entire history, enabling better retention of remote information.
@@ 6 Kalman Filter for Optimally Modelling Fading Memory
In this section, we show how the Kalman Filter (KF) provides a principled solution for constructing an optimal fading memory that accounts for the entire history. We begin by describing the standard Kalman Filter recurrence in the context of memory modeling. However, the KF has a fundamental limitation: its inherently sequential nature makes it impractical for large-scale training on modern hardware accelerators Section˜3.2. To address this, we make simplifying assumptions that makes KF amenable to parallelization on modern hardware accelerators. We then demonstrate that several recent state-space models (DeltaNet, Gated DeltaNet, and Kimi Delta Attention) can be viewed as approximations to the KF recurrence. Specifically, these methods approximate the “optimal" Kalman gain matrix while ignoring dependencies on the past. In contrast, GKA computes the exact Kalman gain by considering the full history. This theoretical advantage translates to improved empirical performance, as we demonstrate in Section˜5.
@@ 6.1 A Dynamical System for Fading Memory
The Kalman filter is a classical algorithm for online optimal inference in linear Gaussian State-Space Models. It gives a principled way to maintain and update a compact state as new noisy observations arrive. The latent state serves as a compressed "memory" of the past. More formally, it is a minimal sufficient statistic that makes past observations conditionally independent of future ones given the state.
@@
Extension to multi-channel measurements. In attention mechanisms, the memory consists of verbatim key-value pairs that can be queried to retrieve past information [Zancato-NeurIPS2024-BMOJO]. Similarly, we want our state to reconstruct past values from their corresponding keys. To achieve this, we extend to a matrix-valued state $S_{t}\in\mathbb{R}^{n\times d}$, where each column independently follows the dynamics in Eq.˜LGM.
@@ 6.2 Kalman Filter for Optimal Inference
where $p$ is the probability density function. The MAP estimate is optimal in the sense that it minimizes the expected squared error between the true state and its estimation given all measurements up to time $t$.
@@ 6.3 Gated KalmaNet: A Steady-State Dynamical System for Large-Scale Training
Collecting all channels, these equations can be written compactly in matrix form as shown in (KF).333with columns of $S_{t}$ transposed to being rows of $S_{t}$ to be consistent with the notation in (KF) and taking the noise variance $r_{t}=\frac{1}{\eta_{t}}$.. A key insight of this work is that the KF recursion for the steady-state model admits an efficient parallel implementation via chunked processing (detailed in Section˜4) that results in Gated KalmaNet.
@@ 6.4 Connection with Existing SSM Layers
for each channel ii. While this added flexibility can improve expressiveness, KDA still assumes $\Sigma_{t}=I_{n}$ and therefore does not optimally consider the entire past when computing its state update. Like DeltaNet and GDN, KDA sacrifices optimality for computational simplicity.
@@ B.1 Forward Pass
Numerical Stability Considerations. Now we analyze the numerical properties of the Chebyshev Iteration. The major computation consists of matrix-vector multiplication; in a batched parallel implementation, this turns out to be matrix-matrix multiplication. For this, the numerical accuracy is well controlled (e.g., in Triton we could specify the accuracy in tl.dot). The update of $\omega_{i}$ in Eq.˜weight schedule might raise numerical concerns as it involves division. That said, we show this division operates in a numerically well-behaved range as $\omega_{i}$ is decreasing with $i$ yet lower bounded by 11:
@@
If $L=\mu$, then $H$ is a scaled identity matrix, and the algorithm is simplified a lot. So we assume $L>\mu$ in what follows. With $L>\mu>0$ we have $\rho\in(0,1)$. Since $\omega_{0}=2$, we have $4-\rho^{2}\omega_{0}\geq 2$ and therefore $0<\omega_{1}\leq 2$. Repeating this argument and we see $\omega_{i}\in(0,2]$ for all ii. By the definition of $\omega_{i}$, to show $\omega_{i}\leq\omega_{i-1}$ is to show
@@
From Lemma˜3 we know that the update of $\omega_{i}$ in Eq.˜weight schedule would not create much numerical concern in a forward pass, as we have $\omega_{i}\in[1,2]$ for all ii. Furthermore, we can bound the rate at which $\omega_{i}$ converges to $\omega_{1}^{*}$:
@@ Appendix C Proof of Lemma˜2
Lemma˜2 describes an interesting phenomenon where, for the CH method (Algorithm˜1), the gradient $d\hat{q}$ obtained from implicit differentiation coincides with the exact gradient dqdq obtained via backpropagation (chain rule). To prove this result, one way is to derive an analytic expression for dqdq (Section˜C.1) and then inspect the recursions. However, this can be algebraically involved. Here, we present a clear proof based on some simple observations.
@@
where the last equality follows, since $H$ is symmetric, which implies $p_{r}(H)$ is symmetric. The proof is finished by observing that $p_{r}(H)\cdot d\xi_{r}$ can be computed via Algorithm˜1 with $H=H,q=d\xi_{r}$ and other parameters, which gives us dqdq.
C.1 The Exact Backward Pass for dqdq and dHdH
Here we show how to obtain the exact gradients of dHdH and dqdq in Algorithm˜1 given the output gradient $d\xi_{r}$, which might be of independent interests. The key insight here is that the Chebyshev iteration can be reversed.
Backward Pass for dq. Let II be the identity matrix of suitable size. To derive a backward pass of Algorithm˜1, we first write down the update of $\xi_{i}$ concisely in the following recursion
@@ C.1 The Exact Backward Pass for dqdq and dHdH
Note that $A_{i}$ is symmetric. Define $d\xi_{i}:=\frac{d\mathcal{L}}{d\xi_{i}}$ for every ii. With some loss function $\mathcal{L}$, assume we are now given $d\xi_{r}$, and our goal is to compute $dq:=\frac{d\mathcal{L}}{dq}$. Since $q$ appears in Eq.˜54 for every $i$, we know $\xi_{0},\xi_{1},\dots,\xi_{r}$ all depend on $q$. Therefore, with $c_{0}:=\frac{2}{L+\mu}$, we have
@@
It remains to compute $d\xi_{i}$ for every ii. Applying the chain rule to Eq.˜54, we obtain
@@ Appendix D Experimental Setup
Training Configurations. All models are trained with the AdamW optimizer with initial learning rate $10^{-3}$, $5\%$ warm-up steps, cosine schedule, gradient clipping with maximum norm 11.
@@
    $k$ = torch.einsum(’...i,...j->...ij’, k, k).cumsum(1)
    $k$ = $k$ / torch.linalg.matrix_norm($k$, ord=’fro’)[..., None, None]
    $k$.diagonal(dim1=-2, dim2=-1).add_(ridge_strength)
@@ Appendix G Effects of Different Regularization Strengths
Synthetic Experiments. First, we generate data as per Fig.˜1a, where the covariance matrix is normalized by its Frobenius norm. In this case we set $\lambda_{t}=a$ for $a$ varying in $\{0.01,0.02,0.05,0.1\}$. Fig.˜8 shows that the maximum regularized residual norm (computed as the maximum of $\|(H_{t}+\lambda_{t}I)\xi_{i}-q\|_{2}$ over all dimensions where $\xi_{i}$ is the estimate of CH at iteration ii) decreases as we enlarge $\lambda_{t}$. This is because having a large $\lambda_{t}$ reduces the condition number. The downside, though, with a large $\lambda_{t}$ is that it reduces the memorization capacity, namely, it might enlarge $\|H_{t}\xi_{i}-q\|_{2}$, the true residual of interest.

### v2

@@ Abstract
As efficient alternatives to softmax Attention, linear State-Space Models (SSMs) achieve constant memory and linear compute, but maintain only a lossy, fading summary of the past, often leading to inferior performance in recall-oriented tasks. We propose Gated KalmaNet (GKA), a layer that accounts for the full past while maintaining SSM-style efficiency. We ground our approach in the Kalman Filter (KF) framework, which provides a principled solution for optimal inference in dynamical systems. We show that several existing SSM layers (DeltaNet, Gated DeltaNet, and Kimi Delta Attention) are approximations to the KF recurrence that assume identity error covariance, thereby ignoring how past measurements (keys and values) should optimally influence state updates. In contrast, GKA computes the exact Kalman gain by maintaining the full error covariance. Under a steady-state assumption that enables parallelization, this reduces to solving an online ridge regression problem with constant memory and linear compute cost. A critical insight is that standard KF equations are numerically unstable in low-precision environments (like bfloat16) and hard to parallelize on modern hardware. We address this through: (1) adaptive regularization with input-dependent gating to control the condition number of the ridge regression for numerical stability, and (2) Chebyshev Iteration, which we show is more stable than conventional iterative solvers in low-precision settings. We further develop hardware-aware chunk-wise kernels to enable efficient training. Empirically, GKA outperforms existing SSM layers (like Mamba2 and Gated DeltaNet) on short-context tasks and achieves more than 10% relative improvement on long-context RAG and LongQA tasks up to 128k tokens.
@@ 2 Prior Work and Preliminaries
Similarly to that of Eq.˜Attn the case with low-rank-plus-identity matrices can often be justified from an optimization perspective. For example, Gated DeltaNet [schlag2021linear, Yang-NeurIPS2024] updates the state via ($I$ is the $D\times D$ identity matrix)
@@
which can be viewed as applying one gradient descent step with step-size $\beta_{t}$ and initialization $\gamma_{t}S_{t-1}$ to the objective
@@ 3.1 Motivation from Kalman Filter
  • In Eq.˜1, the weights of the same residual vary over time as the queries differ at each time, while in Eq.˜3 $i$-th weight $\eta_{i}$ is constant once observed at time $i$. The former results in quadratically many weights—thus a quadratic time complexity—and the latter linearly many.
@@ 4.1 CH with Adaptive Regularization & Weighting
With the weighted covariances $U_{t}:=\sum_{i=1}^{t}\eta_{i}v_{i}k_{i}^{\top}$ and $H_{t}:=\sum_{i=1}^{t}\eta_{i}k_{i}k_{i}^{\top}$, we note that $y_{t}$ can be computed via first solving $(H_{t}+\lambda I)x=q_{t}$ for $x$ and then left-multiplying $U_{t}$. An exact solver (e.g., torch.linalg.solve) can do so with high accuracy by parallelizing over the batch dimension. However, it is inefficient for two reasons. First, it takes $O(D^{3})$ time for every $t$. Second, it requires explicitly forming and materializing all $H_{t}$’s, which entails large I/O costs. In light of this, we resort to first-order iterative methods. These methods use matrix-vector products to enable chunk-wise parallelism over batches without materializing all $H_{t}$’s. Furthermore, they often take $O(D^{2})$ time complexity per iteration and can converge quickly in a few iterations. The iterative method we choose is the Chebyshev Iteration (CH); we proceed to describe its basic idea, with the justification for using CH deferred to Section˜4.2.3.
@@ 4.2.2 Backward Pass
Chunk-wise Recurrence. As indicated, a chunk-wise implementation for computing $A_{i}k_{i}$ is known. On the other hand, computing $B_{i}k_{i}$ is more challenging than $A_{i}k_{i}$, as the additive term $w_{i}H_{i}$ in the backward recursion Eq.˜9 is not necessarily rank-$1$ ; rather, $H_{i}$ itself is defined via the forward recursion in Eq.˜CH. Our contribution here is a derivation for computing $B_{i}k_{i}$ efficiently in a chunk-wise manner.
@@ 4.2.3 Comparison to Other Iterative Solvers
In Fig.˜1b, CG (impl) as a standalone layer has its gradient close to that of the exact solver up to a $10^{-3}$ relative difference. In Fig.˜1c, this difference is amplified to almost $1$ in a 5-layer LLAMA where Eq.˜Attn is replaced with Eq.˜3. This indicates CG (impl) completely deviates from the reference gradient (exact), defeating its purpose of training the network from the regression feedback. In contrast, the gradients of CH (impl) and CH (auto) are eventually close to that of the exact solver either as a single layer (Fig.˜1c) or within multiple layers (Fig.˜1d), up to a $10^{-6}$ difference. Moreover, the curves for CH (impl) and CH (auto) nearly overlap, suggesting that their gradients may be close. The following lemma confirms and formalizes this intuition (see Appendix˜C for a proof), thereby justifying our choice of CH over the alternatives:
@@ 5.1 GKA on Synthetic Associative Recall Tasks
We compare GKA with Attention and other linear SSM baselines on this task. For each memory layer type, we train 2-layer models on MQAR training data and evaluate on a held-out test set. We repeat this experiment for $4$ different learning rates spanning from $10^{-4}$ to $10^{-2}$. As shown in Fig.˜3a, GKA improves upon every other linear SSM baseline at all sequence lengths and model dimensions considered. Note, the complexity of the task increases with increasing sequence length and number of key-value pairs, while larger model dimensions improve memorization capacity through increased state size. The success of our layer can be attributed to our modeling choice: unlike other fading memory designs (like GDN or Mamba2), we construct states based on the optimal MAP estimate conditioned on the entire history, enabling better retention of remote information.
@@ 6 Kalman Filter for Optimally Modelling Fading Memory
In this section, we show how the Kalman Filter (KF) provides a principled solution for constructing an optimal fading memory that accounts for the entire history. We begin by describing the standard Kalman Filter recurrence in the context of memory modeling. However, the KF has a fundamental limitation: its inherently sequential nature makes it impractical for large-scale training on modern hardware accelerators (Section˜3.2). To address this, we make simplifying assumptions that makes KF amenable to parallelization on modern hardware accelerators. We then demonstrate that several recent state-space models (DeltaNet, Gated DeltaNet, and Kimi Delta Attention) can be viewed as approximations to the KF recurrence. Specifically, these methods approximate the “optimal" Kalman gain matrix while ignoring dependencies on the past. In contrast, GKA computes the exact Kalman gain by considering the full history. This theoretical advantage translates to improved empirical performance, as we demonstrate in Section˜5.
@@ 6.1 A Dynamical System for Fading Memory
The Kalman filter is a classical algorithm for online optimal inference in Linear Gaussian State-Space Models. It gives a principled way to maintain and update an estimate of a latent state as new noisy observations arrive. The true latent state serves as a compressed ’memory’ of the past: given the state, the future becomes conditionally independent of past observations.
@@
Extension to multi-channel measurements. In attention mechanisms, the memory consists of verbatim key-value pairs that can be queried to retrieve past information [Zancato-NeurIPS2024-BMOJO]. Similarly, we want our state to reconstruct past values from their corresponding keys. To achieve this, we extend to a matrix-valued state $S_{t}\in\mathbb{R}^{n\times n}$, where each column independently follows the dynamics in Eq.˜LGM.
@@ 6.2 Kalman Filter for Optimal Inference
where $p$ is a probability density function. The MAP estimate is optimal in the sense that it minimizes the expected squared error between the true state and its estimation given all measurements up to time $t$.
@@ 6.3 Gated KalmaNet: A Steady-State Dynamical System for Large-Scale Training
Collecting all channels, these equations can be written compactly in matrix form as shown in (KF)333with columns of $S_{t}$ transposed to being rows of $S_{t}$ to be consistent with the notation in (KF) and taking the noise variance $r_{t}=\frac{1}{\eta_{t}}$.. A key insight of this work is that the KF recursion for the steady-state model admits an efficient parallel implementation via chunked processing (detailed in Section˜4) that results in Gated KalmaNet.
@@ 6.4 Connection with Existing SSM Layers
for each channel $i$. While this added flexibility can improve expressiveness, KDA still assumes $\Sigma_{t}=I_{n}$ and therefore does not optimally consider the entire past when computing its state update. Like DeltaNet and GDN, KDA sacrifices optimality for computational simplicity.
@@ B.1 Forward Pass
Numerical Stability Considerations. Now we analyze the numerical properties of the Chebyshev Iteration. The major computation consists of matrix-vector multiplication; in a batched parallel implementation, this turns out to be matrix-matrix multiplication. For this, the numerical accuracy is well controlled (e.g., in Triton we could specify the accuracy in tl.dot). The update of $\omega_{i}$ in Eq.˜weight schedule might raise numerical concerns as it involves division. That said, we show this division operates in a numerically well-behaved range as $\omega_{i}$ is decreasing with $i$ yet lower bounded by $1$:
@@
If $L=\mu$, then $H$ is a scaled identity matrix, and the algorithm is simplified a lot. So we assume $L>\mu$ in what follows. With $L>\mu>0$ we have $\rho\in(0,1)$. Since $\omega_{0}=2$, we have $4-\rho^{2}\omega_{0}\geq 2$ and therefore $0<\omega_{1}\leq 2$. Repeating this argument and we see $\omega_{i}\in(0,2]$ for all $i$. By the definition of $\omega_{i}$, to show $\omega_{i}\leq\omega_{i-1}$ is to show
@@
From Lemma˜3 we know that the update of $\omega_{i}$ in Eq.˜weight schedule would not create much numerical concern in a forward pass, as we have $\omega_{i}\in[1,2]$ for all $i$. Furthermore, we can bound the rate at which $\omega_{i}$ converges to $\omega_{1}^{*}$:
@@ Appendix C Proof of Lemma˜2
Lemma˜2 describes an interesting phenomenon where, for the CH method (Algorithm˜1), the gradient $d\hat{q}$ obtained from implicit differentiation coincides with the exact gradient $dq$ obtained via backpropagation (chain rule). To prove this result, one way is to derive an analytic expression for $dq$ (Section˜C.1) and then inspect the recursions. However, this can be algebraically involved. Here, we present a clear proof based on some simple observations.
@@
where the last equality follows, since $H$ is symmetric, which implies $p_{r}(H)$ is symmetric. The proof is finished by observing that $p_{r}(H)\cdot d\xi_{r}$ can be computed via Algorithm˜1 with $H=H,q=d\xi_{r}$ and other parameters, which gives us $dq$.
C.1 The Exact Backward Pass for $dq$ and dHdH
Here we show how to obtain the exact gradients of $dH$ and $dq$ in Algorithm˜1 given the output gradient $d\xi_{r}$, which might be of independent interests. The key insight here is that the Chebyshev iteration can be reversed.
Backward Pass for dq. Let $I$ be the identity matrix of suitable size. To derive a backward pass of Algorithm˜1, we first write down the update of $\xi_{i}$ concisely in the following recursion
@@ C.1 The Exact Backward Pass for $dq$ and dHdH
Note that $A_{i}$ is symmetric. Define $d\xi_{i}:=\frac{d\mathcal{L}}{d\xi_{i}}$ for every $i$. With some loss function $\mathcal{L}$, assume we are now given $d\xi_{r}$, and our goal is to compute $dq:=\frac{d\mathcal{L}}{dq}$. Since $q$ appears in Eq.˜54 for every $i$, we know $\xi_{0},\xi_{1},\dots,\xi_{r}$ all depend on $q$. Therefore, with $c_{0}:=\frac{2}{L+\mu}$, we have
@@
It remains to compute $d\xi_{i}$ for every $i$. Applying the chain rule to Eq.˜54, we obtain
@@ Appendix D Experimental Setup
Training Configurations. All models are trained with the AdamW optimizer with initial learning rate $10^{-3}$, $5\%$ warm-up steps, cosine schedule, gradient clipping with maximum norm $1$.
@@
    kk = torch.einsum(’...i,...j->...ij’, k, k).cumsum(1)
    kk = kk / torch.linalg.matrix_norm(kk, ord=’fro’)[..., None, None]
    kk.diagonal(dim1=-2, dim2=-1).add_(ridge_strength)
@@ Appendix G Effects of Different Regularization Strengths
Synthetic Experiments. First, we generate data as per Fig.˜1a, where the covariance matrix is normalized by its Frobenius norm. In this case we set $\lambda_{t}=a$ for $a$ varying in $\{0.01,0.02,0.05,0.1\}$. Fig.˜8 shows that the maximum regularized residual norm (computed as the maximum of $\|(H_{t}+\lambda_{t}I)\xi_{i}-q\|_{2}$ over all dimensions where $\xi_{i}$ is the estimate of CH at iteration $i$) decreases as we enlarge $\lambda_{t}$. This is because having a large $\lambda_{t}$ reduces the condition number. The downside, though, with a large $\lambda_{t}$ is that it reduces the memorization capacity, namely, it might enlarge $\|H_{t}\xi_{i}-q\|_{2}$, the true residual of interest.